HALF −SANDW ICH STRUCTURES OF SCANDIUM AND TITANIUM CYCLOOCTATETRAENE COM-PLEXES

\underline{JUNG SUP LEE}, YUXIU LEI, and DONG-SHENG YANG, Department of Chemistry, University of Kentucky, Lexington, KY 40506-0055.

Unlike benzene, 1,3,5,7-cyclooctatetraene (COT) is not an aromatic molecule and has a tub-like structure in its ground electronic state. However, coordination of Sc or Ti converts the tub-like structure into a planar ring structure. The M-COT (M = Sc, Ti) complexes formed by interactions of laser-vaporized metal atoms and the organic ligand are studied by pulsed field ionization-zero electron kinetic energy photoelectron spectroscopy and density functional theory calculations. The theoretical calculations predict that both Sc/Sc^{+} and Ti/Ti^{+} prefer the η^{8} binding with slightly different M-C distances. The resulting Sc/Sc^{+} and Ti^{+} complexes have C_{4v} molecular symmetry, whereas Ti-COT has C_{2v} symmetry, due to Jahn-Teller distortion. By comparing the experimental measurements and theoretical calculations, the photoelectron spectrum of the Sc complex is assigned to the ^{1}A_{1} ← ^{2}A_{1} transition with a strong 0-0 transition at 42261(5) cm^{−1}. On the other hand, the spectrum of the Ti complex is due to the transitions from two neutral states with similar energies, 3B_{1} and ^{3}B_{2}, to ionic ^{2}A_{1} electronic state. These transitions display strong 0-0 transitions at 40887(5) and 40863(5) cm^{−1}, respectively. In addition, a major vibrational progression (338 and 330 cm^{−1} for the Sc^{+}- and Ti^{+}-COT stretch frequencies, respectively) and a number of vibrational hot bands are observed for the two complexes.